# New Online Communities: Graph Deep Learning on Anonymous Voting Networks to Identify Sybils in Polycentric Governance

### v6

@@ 1   Introduction
A theoretical exploration of endogenous and exogenous models drawn from research on Online Communities, coupled with large scale empirical data and analysis, shapes the foundation of the neoinstitutional theory offered here. Such a theory is necessary to understand the purposes of DAOs and how they operate. In a nod to Elinor Ostrom’s work, we posits that DAOs can effectively govern Digital Common Pool Resources (DCPRs) through polycentric governance, despite their idiosyncratic challenges, if and only if they effectively manage Sybils. Sybils are fraudulent identities used to manipulate voting within digital decentralized governance (similar to ‘sockpuppets’ or ‘non-player characters’ in the system; the term originates in the early 2000s as part of cyber history’s dramatis personae). While traditional governance frameworks readily address the ‘one person, one vote’ principle, the anonymity provided by blockchain technologies complicates the influence of Sybils in decentralized governance.
@@ 2   Deciphering DAOs: Advancing Research in Digital Collective Action
Before describing how the emergence of Digital Common Pool Resources (DCPRs) extends the existing neoinstitutional, resource-management perspective, we offer a brief literature review of the emerging subfield of “DAO Studies” by explicating existing empirical measurements and detailing the conceptual challenges these studies face. Then we briefly connect this perspective to Online Communities literature in Information Systems, in effect linking previous studies on “Web 2.0” to a new phenomenon in “Web3.” Informed by both the dynamics of online communities and commons resource management practices, we gesture towards a theory of digital asset management in a true digital commons. We conclude with an experimental analysis of a large DAO dataset (snapshot.org) by developing a graph deep learning method for detecting Sybils, addressing a key existing challenge to successful decentralized governance.
@@ 3   Emergence and Evolution of DAO Studies
This event, which exposed vulnerabilities of DAOs but also spurred interest in, and the development of, decentralized governance technologies, brought about a significant shift in the discourse surrounding these structures. The 2016 The DAO attack (see [15]) serves as a key event, illuminating emerging issues surrounding the governance and legitimacy of DAOs and contributing to the theoretical foundations of a rapidly emerging DAO Studies.
@@
As the discourse within DAO Studies evolves, we are seeing a shift from insular, participant-derived narratives towards the integration of established socio-economic theories that offer a broader context. A pivotal theoretical framework used to analyze these phenomena has been Hirschman’s paradigm of “Exit, Voice, and Loyalty]] [30]. In this model, ‘Exit’ refers to the option of leaving an organization or community when its quality deteriorates or when disagreements arise, while ‘Voice’ involves expressing dissatisfaction or working towards change from within the organization. ‘Loyalty,’ on the other hand, reflects a commitment to the organization that might inhibit ‘Exit’ and promote ‘Voice.’
@@
In the context of DAOs, ‘Exit’ corresponds to members’ ability to disassociate from the organization, typically by selling their tokens, in response to dissatisfaction or perceived decrease in value or utility. In comparison, ‘Voice’ manifests in the community’s decision-making processes, where token holders can vote on proposals or express their opinions in fora, thereby actively participating in shaping the organization’s trajectory. The traditional notion of ‘Loyalty,’ however, undergoes a transformation in the realm of DAOs. With the advent of ‘Exit to Community’ (E2C) – a concept coined by Schneider and Mannan [52] – ‘Loyalty’ can morph into a mechanism for community ownership. E2C emphasizes that a loyal user or member, rather than simply exiting when dissatisfied, may have the option to help transform the organization into a more community-oriented or user-owned entity, thus introducing a novel form of ‘Loyalty’ that redefines the relationship between community members and the DAO.
@@
Similarly, many scholars have addressed how trust is transformed by DAOs. Much literature on the role of trust in blockchain technology presupposes that its function is to lower transaction costs within organizations, a sentiment summarized by Williamson [60] with his famous theory of Transaction Cost Economics (TCE). Halaburda et al.’s [27] research extends this argument by exploring how blockchain, as a ‘trust-reducing’ technology, can alter the dynamics of transaction governance in organizations, including DAOs. This work explores the intricacies of how blockchain’s inherent features, such as decentralization, transparency, and immutability, contribute to reducing uncertainties and transaction costs that are traditionally managed through hierarchical organizational structures. Halaburda’s insights are particularly relevant in the context of DAOs, which are presupposed to operate on principles of decentralized governance and collective decision-making. By reducing the need for intermediaries and centralized control, blockchain technology enables DAOs to operate with a higher degree of trust among participants (see the ‘Threat Model’ section 5.2.1 for a description of the motivation), fostering a collaborative environment that is less susceptible to opportunistic behaviors, a key concern in TCE.
@@ 3.1   Quantitative and Empirical DAO Research
Goldberg and Schär’s [26] study of Decentraland’s DAO governance analyzes 1,414 governance proposals, involving 45,333 votes from 4,345 addresses, to elucidate the ‘whale effect’ and its influence on governance centralization. Their research revealed critical disparities in voting power and the strategic timing of votes, which impact the democratic integrity of the governance process.
@@
Finally, Bellavitis, Fisch, and Momtaz [5] conduct time series analyses on 2,300 DAOs using a “DAO Analyzer” dataset and studied voting from 1,196 DAOs using a Snapshot dataset. In the “DAO Analyzer” dataset they analyzed the financial returns of investments in DAO tokens, highlighting a key misunderstanding about DAOs and the impact of coin voting. While it is common practice to purchase DAO governance tokens with the expectation that they will yield financial returns like a conventional cryptocurrency, they also are used for governance voting. Coin voting creates a duality of motivations which muddy the conceptual waters, leading us to wonder, when do coins circulate like an economy, return value like an investment, or permit sustainable resource extraction like a commons? While the ‘coin’ in ‘coin voting’ functions like money, we argue below that DAOs are better understood as a commons resource management problem where coin voting determines institutional change through polycentric governance.
@@ 4.2.1 Digital Common Pool Resources
Ostrom’s ideas about polycentric governance can be adopted and adapted to digital assets and their unique governance challenges. Here, we present the novel idea of “Digital Common Pool Resources” (DCPRs) as a digital adaptation of her theory of Common Pool Resources (CPRs). In this context, resources are not physical but digital, like codebases, online communities, or cryptographic assets. Ostrom’s principles find renewed relevance in the rapidly evolving world of blockchain and cryptocurrency, where decentralized, community-based control is paramount. Scholars such as De Filippi and Hassan [24] have further explored how blockchain technologies align with Ostrom’s principles, offering new opportunities for cooperative governance. Ostrom’s neoinstitutional approach provides a foundational framework for understanding self-governance in decentralized digital systems, focusing on the human-centered aspects of these technologies. This approach emphasizes collective action, reciprocity, and shared governance, aligning closely with the core philosophies of decentralization in the digital age.
@@ 5.2.1 Threat Model
The typical free/libre and open source (FLOSS) alternative to decentralized governance (as with DAOs) has distinct challenges when tasked with sustainably and securely managing digital commons. The traditional FLOSS governance models (Linux’s benevolent dictator Linus, Wikipedia’s altruistic editors, Apache’s Foundation model, etc.) all suffer from a reliance on personal trust or legal structuring, neither of which are robust against malicious strategic attack. For the sustainability of digital commons, the decentralization of human decision making is essential, which is why we expect to see more DAOs in the future.
Monero, the leading crime and privacy cryptocurrency, offers an interesting case study of the challenges facing traditional FLOSS management of valuable digital resources. Like many traditional FLOSS software projects, Monero relies on a trusted core team of developers with commit access to a code repository. However, despite years of successful development and increasing use, Monero has recently encountered a potentially fatal organizational issue; a theft of donations (practically all cryptocurrencies encounter such threats eventually). In late 2023 a Monero wallet maintained by several core project members was drained. Accusations of internal theft pointed to a core member (@Luigi) who had recently been arrested (apparently on unrelated charges). This set of events shook up the Monero project which led to a transition to a decentralized governance model. The long term success of Monero hinges on this transition.
In an effort to better understand the threat model facing traditionally-organized (FLOSS) cryptocurrencies like Monero, we used the GitHub API to collect all (2,984) of Monero’s ‘Issues’ and the associated comments. Each GitHub issue with its comments was then processed by a LLM (GPT-4 Turbo), prompted to identify the main issues, relevant actors, sentiment, and to categorize as either governance, security, technical, or financial. The result was further analysed by concatenating the summaries and passing these back into the LLM for further summarizing (to accommodate input token restrictions for the LLM, we vectorized the summaries, trained embeddings with a BERT pre-trained model, and clustered into eight, appropriately sized chunks of input text). Some notable threat issues emerged from this bootstrapped history (see Appendix Appendix: Retrieval Augmented Generation (RAG) of Software Engineering Issues for Monero Cryptocurrency: Hierarchical GPT-4 Turbo for Automated Histories*), including debates about infrastructure changes, community engagement and moderation, and finally, two security issues that reflect broader issues of transparency and governance. This hierarchical Retrieval Augmented Generation (RAG) method can be used at scale for rapidly generating governance histories.
@@ 5.2.2 Defining Sybils and Deanonymizing Social Networks
Similarly, more recent work has been successful by using an open source information (OSI) approach to identify additional auxiliary graphs for deanonymizing Bitcoin users [8]. This research focuses on the ability to walk the transaction graph and uses a ‘follow the money’ approach to track down auxiliary information sources. However, in practical terms, this approach relies on poor security behaviour (high k-anonymity) and investigations often lead to an anonymizing coin mixer or a regulated exchange that can only be further investigated by legal authorities.
@@
The most effective research on deanonymizing Bitcoin transactions uses real labelled datasets and intuitive machine learning algorithms (Akcora et al. [1] and Paquet-Clouston et al. [45]) or graph deep learning (Zhou et al. [65]), but these studies still have significant limitations. Paquet-Clouston et al. [45] attempt to trace ransomware payments by constructing graphs from 7222 labelled seeds originating from a convenience sample of known ransomware addresses. They construct an address graph and a cluster graph, the latter uses the logic of Sybils to combine nodes that have shared inputs (multiple-input clustering heuristics). However, their analysis is purely descriptive and does not attempt to make predictions. Akcora et al. [1] also investigate ransomware transactions, use convenience sample of three datasets (which includes the Paquet-Clouston et al. dataset), and construct a transaction graph and an address graph. They use multiple heuristics, develop graph features, use DBSCAN and k-means algorithms to cluster addresses and a naive similarity search.
@@
Finally, Zhou et al. [65] attempt Ethereum address clustering, or what we would call ‘Sybil identification.’ They construct a trimmed graph of Ethereum transactions, develop node features, and train a graph neural network (GNN) to search subgraph node embeddings using a hierarchical (weighted) graph attention autoencoder. As is typical, they use several pooling layers and non-linear activation layers (LeakyReLu). They use cross-entropy loss for the classifier. They improve the performance of the default GNN by using contrastive subgraph learning. This self-supervised methodology, while in a different context, is similar to the approach detailed below. The significant difference is that they classified nodes directly using the GNN, whereas we use develop the embeddings first and then classify using a embeddings search, an encoder-decoder model. The second step is necessary to detect Sybils because Sybils are by definition unknown and with limited connectivity to other voter nodes.
@@
In the context of decentralized governance, however, Sybils are identified using our more relaxed definition that circumvents k-anonymity. In this context, our method for identifying Sybils is a weak form of deanonymization but one that reveals no personally identifiable information and unlike most previous methods does not require auxiliary or labelled graphs (although it can benefit from such information). Supervised techniques for deanonymization will not effectively identify Sybils because, in a social graph characterized by strong anonymity, Sybils may create a single, disconnected edge and thus lack sufficient connectivity to be identified by walk-based deanonymization approaches. Naive representations of Sybils contain only 1 bit of structural information. Indeed, we assume that a majority of Sybils are singletons. The existence of a labelled subgraph of known voters, linked to Ethereum Name Service (ENS) accounts, provides little help in identifying isolated Sybils, for the same reason (but can be used to validate clusters).
@@ 5.3   Experimental Sybil Identification Methodology
The dataset covers governance voting for roughly 35,000 DAOs between 2020-07-18 and 2023-02-27 (959 days, 2.6 years), which covers approximately 7.2 million voting transactions. However, naive counts of DAOs, voters, and their governance proposals are ineffective due to noise. For instance, upon manual inspection the dataset reveals many registered DAOs that failed to launch or were clearly set up as tests, and many governance proposals have durations far too short to be practical (see Figure 8).
@@
Even in the ‘simplest’ case of Bitcoin, privacy preserving blockchains present a number of obstacles to forensic, especially ‘taint’ analysis. Rather straightforwardly, since Bitcoin uses an unspent transaction output (UTXO) accounting model, it is possible to determine when wallet addresses that require the exact same knowledge to spend can be considered equivalent (BlockSci’s “equivalent addresses”), which is a type of direct Sybil identification. However, even when assuming users are not following privacy-maximizing address rotation practices (about half of Bitcoin transactions), tracing financial flows still requires linking inputs and outputs, addressing “change” wallets, and surmounting coinJoin transactions and coin mixing services.

### v7

@@ 1   Introduction
A theoretical exploration of endogenous and exogenous models drawn from research on Online Communities, coupled with large scale empirical data and analysis, shapes the foundation of the neoinstitutional theory offered here. Such a theory is necessary to understand the purposes of DAOs and how they operate. In a nod to Elinor Ostrom’s work, we posit that DAOs can effectively govern Digital Common Pool Resources (DCPRs) through polycentric governance, despite their idiosyncratic challenges, if and only if they effectively manage Sybils. Sybils are fraudulent identities used to manipulate voting within digital decentralized governance (similar to ‘sockpuppets’ or ‘non-player characters’ in the system; the term originates in the early 2000s as part of cyber history’s dramatis personae). While traditional governance frameworks readily address the ‘one person, one vote’ principle, the anonymity provided by blockchain technologies complicates the influence of Sybils in decentralized governance.
@@ 2   Deciphering DAOs: Advancing Research in Digital Collective Action
Before describing how the emergence of Digital Common Pool Resources (DCPRs) extends the existing neoinstitutional, resource-management perspective, we offer a brief literature review of the emerging subfield of “DAO Studies” by explicating existing empirical measurements and detailing the conceptual challenges these studies face. Then we briefly connect this perspective to Online Communities literature in Information Systems, in effect linking previous studies on “Web 2.0” to a new phenomenon in “Web3.” Informed by both the dynamics of online communities and commons resource management practices, we gesture towards a theory of digital asset management in a true digital commons. We conclude with an experimental analysis of a large DAO dataset (snapshot.org) and develop a graph deep learning method for detecting Sybils, addressing a key existing challenge to successful decentralized governance.
@@ 3   Emergence and Evolution of DAO Studies
This event, which exposed vulnerabilities of DAOs but also spurred interest in and the development of decentralized governance technologies, brought about a significant shift in the discourse surrounding these structures. The 2016 The DAO attack (see [16]) serves as a key event, illuminating emerging issues surrounding the governance and legitimacy of DAOs and contributing to the theoretical foundations of a rapidly emerging DAO Studies.
@@
As the discourse within DAO Studies evolves, we are seeing a shift from insular, participant-derived narratives towards the integration of established socio-economic theories that offer a broader context. A pivotal theoretical framework used to analyze these phenomena has been Hirschman’s paradigm of “Exit, Voice, and Loyalty” [31]. In this model, ‘Exit’ refers to the option of leaving an organization or community when its quality deteriorates or when disagreements arise, while ‘Voice’ involves expressing dissatisfaction or working towards change from within the organization. ‘Loyalty,’ on the other hand, reflects a commitment to the organization that might inhibit ‘Exit’ and promote ‘Voice.’
@@
In the context of DAOs, ‘Exit’ corresponds to members’ ability to disassociate from the organization, typically by selling their tokens, in response to dissatisfaction or perceived decrease in value or utility. In comparison, ‘Voice’ manifests in the community’s decision-making processes, where token holders can vote on proposals or express their opinions in online fora, thereby actively participating in shaping the organization’s trajectory. The traditional notion of ‘Loyalty,’ however, undergoes a transformation in the realm of DAOs. With the advent of ‘Exit to Community’ (E2C) – a concept coined by Schneider and Mannan [53] – ‘Loyalty’ can morph into a mechanism for community ownership. E2C emphasizes that a loyal user or member, rather than simply exiting when dissatisfied, may have the option to help transform the organization into a more community-oriented or user-owned entity, thus introducing a novel form of ‘Loyalty’ that redefines the relationship between community members and the DAO.
@@
Similarly, many scholars have addressed how trust is transformed by DAOs. Much literature on the role of trust in blockchain technology presupposes that its function is to lower transaction costs within organizations, a sentiment summarized by Williamson [61] with his famous theory of Transaction Cost Economics (TCE). Halaburda et al.’s [28] research extends this argument by exploring how blockchain, as a ‘trust-reducing’ technology, can alter the dynamics of transactional governance in organizations, including DAOs. This work explores the intricacies of how blockchain’s inherent features, such as decentralization, transparency, and immutability, contribute to reducing uncertainties and transaction costs that are traditionally managed through hierarchical organizational structures. Halaburda’s insights are particularly relevant in the context of DAOs, which are presupposed to operate on principles of decentralized governance and collective decision-making. By reducing the need for intermediaries and centralized control, blockchain technology enables DAOs to operate with a higher degree of trust among participants (see the ‘Threat Model’ section 5.2.1 for a description of the motivation), fostering a collaborative environment that is less susceptible to opportunistic behaviors, a key concern in TCE.
@@ 3.1   Quantitative and Empirical DAO Research
Goldberg and Schär’s [27] study of Decentraland’s DAO governance analyzes 1,414 governance proposals, involving 45,333 votes from 4,345 addresses to elucidate the ‘whale effect’ and its influence on governance centralization. Their research revealed critical disparities in voting power and the strategic timing of votes, which impact the democratic integrity of the governance process.
@@
Finally, Bellavitis, Fisch, and Momtaz [5] conduct time series analyses on 2,300 DAOs using a “DAO Analyzer” dataset and studied voting from 1,196 DAOs using a Snapshot dataset. In the “DAO Analyzer” dataset they analyzed the financial returns of investments in DAO tokens, highlighting a key misunderstanding about DAOs and the impact of coin voting. While it is common practice to purchase DAO governance tokens with the expectation that they will yield financial returns like a conventional cryptocurrency, they are alsio used for governance voting. Coin voting creates a duality of motivations which muddy the conceptual waters, leading us to wonder, when do coins circulate like an economy, return value like an investment, or permit sustainable resource extraction like a commons? While the ‘coin’ in ‘coin voting’ functions like money, we argue below that DAOs are better understood as a commons resource management problem where coin voting determines institutional change through polycentric governance.
@@ 4.2.1 Digital Common Pool Resources
Ostrom’s ideas about polycentric governance can be adopted and adapted to digital assets and their unique governance challenges. Here, we present the novel idea of “Digital Common Pool Resources” (DCPRs) as a digital adaptation of her theory of Common Pool Resources (CPRs). In this context, resources are not physical but digital, like codebases, online identities, or cryptographic assets. Ostrom’s principles find renewed relevance in the rapidly evolving world of blockchain and cryptocurrency, where decentralized, community-based control is paramount. Scholars such as De Filippi and Hassan [25] have further explored how blockchain technologies align with Ostrom’s principles, offering new opportunities for cooperative governance. Ostrom’s neoinstitutional approach provides a foundational framework for understanding self-governance in decentralized digital systems, focusing on the human-centered aspects of these technologies. This approach emphasizes collective action, reciprocity, and shared governance, aligning closely with the core philosophies of decentralization in the digital age.
@@ 5.2.1 Threat Model
The typical free/libre and open source (FLOSS) alternative to decentralized governance has distinct challenges when tasked with sustainably and securely managing digital commons. The traditional FLOSS governance models (Linux’s benevolent dictator Linus, Wikipedia’s altruistic editors, Apache’s Foundation model, etc.) all suffer from a reliance on personal trust or legal structuring, neither of which are robust against malicious strategic attack or free to develop in arbitrary (perhaps anti-social or illegal) ways. For the sustainability of digital commons, the decentralization of human decision making is essential, which is why we expect to see more DAOs in the future.
Monero, the leading crime and privacy cryptocurrency, offers an interesting case study of the challenges facing traditional FLOSS management of valuable digital resources. Like many traditional FLOSS software projects, Monero relies on a trusted core team of developers with commit access to a code repository. However, despite years of successful development and increasing use, Monero has recently encountered a potentially fatal organizational issue; a theft of donations (practically all cryptocurrencies encounter such threats eventually). In late 2023 a Monero wallet maintained by several core project members was drained. Accusations of internal theft pointed to a core member (@Luigi) who had recently been arrested (apparently on unrelated charges). This set of events shook up the Monero project which has led to efforts towards transition to a decentralized governance model. The long term success of Monero hinges on this transition.
In an effort to better understand the threat model facing traditionally-organized (FLOSS) cryptocurrencies like Monero, we used the GitHub API to collect all (2,984) of Monero’s ‘Issues’ and the associated comments. Each GitHub issue with its comments was then processed by a LLM (GPT-4 Turbo), prompted to identify the main issues, relevant actors, sentiment, and to categorize as either governance, security, technical, or financial. The result was further analysed by concatenating the summaries and passing these back into the LLM for further summarizing (to accommodate input token restrictions for the LLM, we vectorized the summaries, trained embeddings with a BERT pre-trained model, and clustered into appropriately sized chunks of input text). Some notable threat issues emerged from this bootstrapped history (see Appendix Appendix: Retrieval Augmented Generation (RAG) of Software Engineering Issues for Monero Cryptocurrency: Hierarchical GPT-4 Turbo for Automated Histories*), including debates about infrastructure changes, community engagement and moderation, and finally, two security issues that reflect broader issues of transparency and governance. This hierarchical Retrieval Augmented Generation (RAG) method can be used at scale for rapidly generating governance histories.
@@ 5.2.2 Defining Sybils and Deanonymizing Social Networks
Similarly, more recent work has successfully used an open source information (OSI) approach to identify additional auxiliary graphs for deanonymizing Bitcoin users [8]. This research focuses on the ability to walk the transaction graph and uses a ‘follow the money’ approach to track down auxiliary information sources. However, in practical terms, this approach relies on poor security behaviour (high k-anonymity) and investigations often lead to an anonymizing coin mixer or a regulated exchange that can only be further investigated by legal authorities.
@@
The most effective research on deanonymizing Bitcoin transactions uses real labelled datasets and intuitive machine learning algorithms (Akcora et al. [1] and Paquet-Clouston et al. [46]) or graph deep learning (Zhou et al. [66]), but these studies still have significant limitations. Paquet-Clouston et al. [46] attempt to trace ransomware payments by constructing graphs from 7222 labelled seeds originating from a convenience sample of known ransomware addresses. They construct an address graph and a cluster graph; the latter uses the logic of Sybils to combine nodes that have shared inputs (multiple-input clustering heuristics). However, their analysis is purely descriptive and does not attempt to make predictions. Akcora et al. [1] also investigate ransomware transactions, use convenience sample of three datasets (which includes the Paquet-Clouston et al. dataset), and construct a transaction graph and an address graph. They use multiple heuristics, develop graph features, use density-based clustering (DBSCAN) and k-means algorithms to cluster addresses with a similarity search.
@@
Finally, Zhou et al. [66] attempt Ethereum address clustering, or what we would call ‘Sybil identification.’ They construct a trimmed graph of Ethereum transactions, develop node features, and train a graph neural network (GNN) to search subgraph node embeddings using a hierarchical (weighted) graph attention autoencoder. As is typical, they use several pooling layers and non-linear activation layers (LeakyReLu). They use cross-entropy loss for the classifier. They improve the performance of the default GNN by using contrastive subgraph learning. This self-supervised methodology, while in a different context, is similar to our approach detailed below. The significant difference is that they classified nodes directly using the GNN, whereas we use develop the embeddings first and then classify using a embeddings search, an encoder-decoder model. The second step is necessary to detect Sybils because Sybils are by definition unknown and with limited connectivity to other voter nodes.
@@
In the context of decentralized governance, however, Sybils are identified using our more relaxed definition that circumvents k-anonymity. In this context, our method for identifying Sybils is a weak form of deanonymization but one that reveals little to no personally identifiable information and unlike most previous methods does not require auxiliary or labelled graphs (although it can benefit from such information). Supervised techniques for deanonymization will not effectively identify Sybils because, in a social graph characterized by strong anonymity, Sybils may create a single, disconnected edge and thus lack sufficient connectivity to be identified by walk-based deanonymization approaches. Naive representations of Sybils contain only 1 bit of structural information. Indeed, we assume that a majority of Sybils are singletons. The existence of a labelled subgraph of known voters, linked to Ethereum Name Service (ENS) accounts, provides little help in identifying isolated Sybils, for the same reason (but can be used to help validate clusters).
@@ 5.3   Experimental Sybil Identification Methodology
Our dataset covers governance voting for roughly 35,000 DAOs between 2020-07-18 and 2023-02-27 (959 days or 2.6 years), which covers approximately 7.2 million voting transactions. However, naive counts of DAOs, voters, and their governance proposals are ineffective due to noise. For instance, upon manual inspection the dataset reveals many registered DAOs that failed to launch or were clearly set up as tests, and many governance proposals have durations far too short to be practical (see Figure 8).
@@
Even in the ‘simplest’ case of Bitcoin, privacy preserving blockchains present a number of obstacles to forensic and especially taint analysis. Rather straightforwardly, since Bitcoin uses an unspent transaction output (UTXO) accounting model, it is possible to determine when wallet addresses that require the exact same knowledge to spend can be considered equivalent (BlockSci’s “equivalent addresses”), which is a type of direct Sybil identification. However, even when assuming users are not following privacy-maximizing address rotation practices (about half of Bitcoin transactions), tracing financial flows still requires linking inputs and outputs, addressing “change” wallets, and surmounting coinJoin transactions and coin mixing services.